OpenStudy (anonymous):

8 to what power = 16???

OpenStudy (anonymous):

It is: 84/3 = 16

OpenStudy (elisaneedshelp):

Yea i dont think 8 to the power of anything could be 16

OpenStudy (raden):

let it is x so, we have an equation : 8^x = 16 setting that 8 = 2^3 and 16 = 2^4, get 8^x = 16 (2^3)^x = 2^4 2^(3x) = 2^4 now the bases are same, just take the power of both and solve for x : 3x = 4 x = 4/3 so, it is 4/3

OpenStudy (elisaneedshelp):

i stand corrected

OpenStudy (elisaneedshelp):

yet again math isnt my best

OpenStudy (anonymous):

Thank youuuuu :)

OpenStudy (raden):

you're welcome